### v1

# Expansion Span: Combining Fading Memory and Retrieval in Hybrid State Space Models

Abstract
The ‘state’ of State Space Models (SSMs) represents their memory, which fades exponentially over an unbounded span. By contrast, Attention-based models have ‘eidetic’ (i.e., verbatim, or photographic) memory over a finite span (context size). Hybrid architectures combine State Space layers with Attention, but still cannot recall the distant past and can access only the most recent tokens eidetically. Unlike current methods of combining SSM and Attention layers, we allow the state to be allocated based on relevancy rather than recency. In this way, for every new set of query tokens, our models can ‘eidetically’ access tokens from beyond the Attention span of current Hybrid SSMs without requiring extra hardware resources. We describe a method to expand the memory span of the hybrid state by ‘reserving’ a fraction of the Attention context for tokens retrieved from arbitrarily distant in the past, thus expanding the eidetic memory span of the overall state. We call this reserved fraction of tokens the ‘expansion span,’ and the mechanism to retrieve and aggregate it ‘Span-Expanded Attention’ (SE-Attn). To adapt Hybrid models to using SE-Attn, we propose a novel fine-tuning method that extends LoRA to Hybrid models (HyLoRA) and allows efficient adaptation on long spans of tokens. We show that SE-Attn enables us to efficiently adapt pre-trained Hybrid models on sequences of tokens up to 8 times longer than the ones used for pre-training. We show that HyLoRA with SE-Attn is cheaper and more performant than alternatives like LongLoRA when applied to Hybrid models on natural language benchmarks with long-range dependencies, such as PG-19, RULER, and other common natural language downstream tasks.
1 Introduction
State Space Models are able to process sequences with an unbounded number of tokens by maintaining a fixed-size state. However, this state is lossy and information about early tokens ‘fades’ as more inputs are processed. In contrast, Transformer models have a state determined by the number of tokens in their input sequence and are able to access information from all past tokens in their context ‘eidetically.’ However, they do so at the cost of extra compute and memory. Recent Hybrid models [38, 12, 8] augment SSMs with Attention layers in an effort to counteract SSMs’ ‘fading’ memory. However, Attention layers only aggregate information by recency (i.e., they process the keys and values of the most recent tokens up to hardware limitations). Hence, any token that lies beyond the Attention’s limited span can only be approximately recalled through the SSMs’ state. This limits the effectiveness of Hybrid SSMs when applied to long sequences of tokens, especially when compute and memory resources are limited. In fact, while recent pre-trained Hybrid models have demonstrated strong performance on common downstream and recall-intensive language tasks, they are trained following Transformer’s pre-training practice: they are either trained on a fixed, relatively short window of tokens when compute is limited—which often is set to 2048 as in LLaMA1 [31], or 4096 as in LLaMA2 [32]—or are progressively fine-tuned on longer sequences after the pre-training stage [22, 12, 9].
In this work, we modify Hybrid SSMs’ Attention layers to allow their state to be allocated by relevancy rather than recency, allowing our models to retrieve information that would otherwise fall outside their Attention span. To do so, we propose Span-Expanded Attention (SE-Attn), a selection mechanism that complements the ‘fading’ memory of SSMs with ‘eidetically’ retrieved tokens.
Our method allows long context fine-tuning of pre-trained Hybrid models on sequences longer than the ones used for pre-training without significant compute/memory overhead. While Transformer-based fine-tuning strategies for processing long sequences can also be applied to Hybrid models, they typically work under the assumption that information in the Attention layers is aggregated by recency and are thus more expensive as sequences get longer. For example, extending a LLaMA model from a 2k to 8k context size can require up to 32 A100 GPUs [2]. While there exist efficient LoRA-based methods [17] to bring this cost down, we show that Transformer-based methods tend to not work as well for Hybrid models as they do for Transformers. We show that this is mostly due to the fact that combining LoRA (on Attention layers) with SSM layers is not expressive enough to model long-range dependencies beyond the pre-training context size.
To overcome such limitations, we propose to modify the standard LoRA recipe to fit the Hybrid model structure better. Namely, we employ a fine-tuning strategy similar to LoRA+ [3] on Attention layers, while allowing the SSMs’ recurrent parameters to be adapted as well. In particular, building on previous observations [38, 37], we also adapt the 1D convolutional layers, which we empirically validate provide the best results at a reduced computational cost.
We make the following contributions:
  1. 1.
    We propose SE-Attn, a novel selection mechanism that ‘reserves’ a fraction of the context of Hybrid SSMs’ Attention layers for tokens that are retrieved from the past based on their relevance to the current query.
  2. 2.
    We empirically show that Transformer-based methods for long-context fine-tuning are sub-optimal when applied to Hybrid SSMs. Therefore, we propose HyLoRA, an extension of LoRA+ [3] for efficient and effective fine-tuning of Hybrid SSMs.
  3. 3.
    We show that combining SE-Attn and HyLoRA on Hybrid models reliably extends their context size up to 8×8$\times$8 × their pre-training size. Further, we show improved performance over existing adaptation methods on common natural language tasks in the LM Harness Evaluation suite [11] and long context tasks in RULER [16].
2 Background and Related Work
Attention on long contexts. Since the introduction of Self-Attention [33], a significant amount of research has been made to reduce its quadratic cost in the sequence length at training time. To achieve this, state of the art methods usually approximate the Attention mechanism with sparse/linearized versions. For example, Reformer [21] uses locality-sensitive-hashing to group tokens with similar embeddings, allowing the model to only attend to a subset of tokens rather than the entire sequence. Longformer [1] combines local Attention, whereby each token is restricted to attending only to a window of neighboring tokens, and global Attention, whereby a few tokens can attend to all tokens. Meanwhile, Linformer [35] uses projection matrices to project keys and values into a lower-dimensional space along the sequence dimension.
Different from these approximate Attention methods, other works have proposed to endow Transformers models with ‘compressed’ memory tokens that are updated dynamically and causally over sliding windows on entire sequence chunks. For example, Transformer-XL [4] and Infini-Attention [24] segment an input sequence into chunks and process them sequentially while maintaining a complementary set of tokens whose purpose is to summarize the older ones. In contrast to these works, our Attention retrieves relevant tokens ‘eidetically,’ i.e., we retrieve tokens rather than maintain a compressed representation of the past. Most similar to our method is Landmark Attention [23], which inserts landmark tokens into the input at fixed block intervals and trains these tokens to act as summaries of their corresponding blocks via a grouped softmax attention; these summary tokens are then used to index and retrieve relevant input tokens when processing future segments. Our Attention aims to integrate retrieval natively, without the need for external landmark tokens or complex Softmax procedures.
State Space Models. While a great effort has been made to improve the efficiency of Transformer models, a recent line of work has explored efficient alternative ‘linear’ architectures. In particular, State Space Models (SSMs) [13, 15, 14, 36, 30] have emerged as promising competitors to Transformer models due to their efficient scaling and strong empirical performance. Numerous variations of SSMs have been proposed, some closely resembling Linear Attention [30] or Linear-Time Invariant dynamical systems [15], while others introduce novel adaptive/gated state updates [36, 13, 6, 25]. Despite their differences, all follow the same basic working principle that is inspired by classical state space models [19]: they process the input sequence by maintaining a fixed-size state which acts as a compressed (lossy) representation of all the processed tokens. However, when implemented in hardware, the state must have finite precision and therefore ‘fades’ as more samples are processed. To overcome this limitation, the most successful SSMs are typically hardware-aware implementations that efficiently utilize modern GPUs/TPUs. These implementations use highly parallelizable and scalable primitives, such as associative scans [13, 8], chunking mechanisms [6, 36], and techniques that avoid the materialization of the entire state in slow high bandwidth memory [13].
Hybrid State Space Models. While extending the recurrent state in SSMs layers has led to performant models, they typically fall short on tasks that require recalling information from the distant past [34, 18]. Hybrid State Space Models have been introduced to solve this limitation and are typically designed to complement the SSMs’ ‘fading’ state with Attention layers [7, 8, 22, 12]. While early architectures simply stack SSMs and Attention layers with different blending ratios [34, 13, 6] or replace full Attention layers with Sliding Window Attention [8], more performant and sophisticated hybrid architectures have been recently proposed [12, 38]. In particular, B’MOJO [38] complements SSMs’ fading state with a form of ‘eidetic’ memory whereby SSM layers are mixed with Sliding Window Attention and tokens in the window can attend to a set of tokens from the past that are deemed difficult to predict using an asynchronous causal selection mechanism. Similarly, in our work, we mix SSM layers with an Attention layer that operates over chunks of tokens, but we enable these chunks to dynamically retrieve tokens from the past that are most relevant to the chunk’s tokens.
3 Span-Expanded Attention
Figure 1: Span-Expanded Attention (SE-Attn) overview. SE-Attn is a sparse Attention mechanism used to expand the memory span of Hybrid SSMs. We do so by reserving a fraction of the Attention context for tokens retrieved arbitrarily far back in the past, and use summary tokens to efficiently look back with a reduced compute cost. We call this reserve the ‘expansion span,’ and we populate it with blocks of previous tokens (memory blocks). When new tokens arrive, a similarity-based search compares the queries with past memory blocks to decide which memory blocks are most relevant. Then, these retrieved memory blocks are jointly processed with the queries via Attention. While the final Attention mechanism always processes a fixed number of tokens, it can have a longer span since tokens from arbitrarily far back in the past can be retrieved.
Our goal is to enable pre-trained Hybrid SSMs to accurately process sequences with a larger number of tokens than were used for pre-training. For a sequence of $L$ tokens with model dimension $d_{\textrm{model}}$, the computational complexity of standard Self-Attention is $O(d_{\textrm{model}}L^{2})$. For very large $L$, computing Self-Attention can be prohibitively expensive. Training models with a large $L$ is particularly challenging, as training has the additional cost of computing gradients, further limiting memory resources. To address this, we propose Span-Expanded Attention (SE-Attn), a drop-in replacement for standard Self-Attention in Hybrid SSMs. To train SE-Attn, we propose HyLoRA, a variant of LoRA+ [3] specifically tailored to Hybrid models. A schematic of our SE-Attn is provided in Figure 1.
3.1 Attention
At the heart of modern LLMs is the Attention mechanism, whose role is to construct a contextualized representation of their input. The input to Attention, $x\in\mathbb{R}^{L\times d}$, is a tensor of $L$ tokens, each with dimension $d$. Attention is parameterized by $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{\textrm{model}}}$ and $W_{o}\in\mathbb{R}^{d_{\textrm{model}}\times d}$, which are used to construct linear projections of the input: $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}\in\mathbb{R}^{L\times d_{\textrm{model}}}$. After adding positional information to the keys and queries, the output of the Attention layer is [^1] For simplicity, we consider Single-Head Attention in this exposition. $Attention(Q,K,V)=\left[\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{d_{\textrm{model}}}}\right)V\right]W_{o}\in\mathbb{R}^{L\times d}$. The realization of the attention score matrix, $QK^{T}\in\mathbb{R}^{L\times L}$, grows quadratically in the sequence length, $L$. Moreover, the output of Attention is typically followed by a feed-forward layer (FFN), which expands and contracts the dimension of the input. While these FFNs are generally regarded as being the primary computational bottlenecks of Attention-based models, when the sequence length exceeds the expanded dimension, Attention becomes the primary bottleneck. Since LLMs consist of many layers of Attention, this computation is particularly expensive during training, where activations are cached. To make the Attention mechanism computationally amenable to longer contexts during training, we draw inspiration from RAG and Sparse Attention methods.
3.2 Amnesic Attention
The crux of our method is in the chunking of the $L$ tokens in the input sequence $x$, into chunks of size $M$. Namely, we begin by computing projections $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}$ and adding positional information as in standard Attention. Next, each of the $Q,K,V\in\mathbb{R}^{L\times d_{\textrm{model}}}$ projections are split into $T=\frac{L}{M}$ chunks of $M$ tokens in each along the sequence dimension, yielding $Q_{i},K_{i},V_{i}\in\mathbb{R}^{M\times d_{\textrm{model}}}$ for $i=1,\ldots,T$. A naive way to reduce the quadratic cost of Attention is by applying it independently on each of these chunks, $A_{i}=Attention(Q_{i},K_{i},V_{i})$ and then concatenate and project them to get the final output, $\textrm{Concatenate}(A_{1},A_{2},\ldots,A_{T})$. However, since the chunks are processed independently, there is no information exchanged between them. Hence, this naive—though efficient—Attention mechanism, which we refer to as “SE-Attn-NoMem”, cannot model time dependencies on contexts larger than the chunk size.
3.3 Eidetic Retrieval Attention
In this section, we improve upon SE-Attn-NoMem by allowing different chunks to exchange information while minimizing compute; we call this Attention mechanism SE-Attn. To this end, we augment the processing of each chunk with a mechanism to retrieve tokens from previous chunks. In particular, we allow chunk $i$ to retrieve tokens from chunks $11,2,\ldots,i-$ 1 and, when the most relevant chunks are selected, append their tokens to its context (its “expansion span”). SE-Attn takes as input a sequence, $x\in\mathbb{R}^{L\times d}$ and computes projections $Q=xW_{Q}$, $K=xW_{K}$, $V=xW_{V}$ followed by the addition of RoPE [29] embeddings as in standard Attention. As described in Sec. 3.2, $Q,K,V$ are split into $T$ chunks with $M$ tokens in each, yielding tuples $(Q_{i},K_{i},V_{i})$, $i=1,\ldots,T$. Additionally, $Q,K,V$ are split into a second set of $U$ chunks with $S$ tokens in each, yielding tuples $(Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j})$ where $Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j}\in\mathbb{R}^{S\times d_{\textrm{model}}}$ for $j=1,\ldots,U$. We refer to these tuples as “memory blocks”; in SE-Attn, the query from each chunk, $Q_{i}$, attends not only to $K_{i}$, but also to a set of retrieved memory blocks which populate SE-Attn’s expansion span. In particular, each $Q_{i}$ retrieves $k$ (top-$k$) key/value memory blocks from the past [^2] By “the past,” we mean tokens in the sequence that came before tokens in the chunk.: $\left(K^{\textrm{Mem}}_{\phi_{i}(U)_{1}},V^{\textrm{Mem}}_{\phi_{i}(U)_{1}}\right),\ldots,\left(K^{\textrm{Mem}}_{\phi_{i}(U)_{k}},V^{\textrm{Mem}}_{\phi_{i}(U)_{k}}\right)$ where $\phi_{i}(U)_{j}$ denotes the index of the $j$-th memory block selected by the $i$-th chunk. These retrieved blocks are then appended to the chunks’ keys and values, and SE-Attn computes the Attention output for the $i$th chunk as in Equation 3.
$\tilde{K}$ $=\textrm{Concatenate}(K^{\textrm{Mem}}_{\phi_{i}(U)_{1}},\ldots,K^{\textrm{Mem}}_{\phi_{i}(U)_{k}},K_{i})$ (1)
$\tilde{V}$ $=\textrm{Concatenate}(V^{\textrm{Mem}}_{\phi_{i}(U)_{1}},\ldots,V^{\textrm{Mem}}_{\phi_{i}(U)_{k}},V_{i})$ (2)
$A_{i}^{\textrm{SE-Attn}}$ $=Attention(Q_{i},\tilde{K}_{i},\tilde{V}_{i})$ (3)
Afterwards, each chunk’s Attention outputs are concatenated and projected to obtain the SE-Attn layer output:
$o^{\textrm{SE-Attn}}$ $=\textrm{Concatenate}(A_{1}^{\textrm{SE-Attn}},A_{2}^{\textrm{SE-Attn}},\dots,A_{T}^{\textrm{SE-Attn}}).$ (4)
Note that Equation 3 is a form of Cross-Attention since we are not concatenating memory query tokens to the chunk’s query. This is to preserve causality, as the retrieved tokens cannot attend to the chunk’s tokens. Next, we discuss how the memory tokens are retrieved.
Memory Retrieval. Each chunk $x_{i}$ must judiciously select which memory blocks to retrieve from the past. To do this efficiently, we associate a 1-dimensional tensor, $c_{j}\in\mathbb{R}^{d_{\textrm{model}}}$, to each of the $j=1,\ldots,U$ memory blocks which act as a compressed representation of each memory block. To determine which memory blocks $Q_{i}$ should attend to, we compute a “relevancy score” between $Q_{i}$ and each $c_{j}$, which measures how relevant memory block $j$ is to chunk $i$. This relevancy score is implemented with a Cross-Attention score between chunks and compressed memory block representations. More specifically, recalling that $Q_{i}\in\mathbb{R}^{M\times d_{\textrm{model}}}$, we compute relevancy score $R_{ij}\in\mathbb{R}$ between chunk $i\in\{1,2,\ldots,T\}$ and memory block $j\in\{1,2,\ldots,U\}$ as follows:
$R_{ij}$ $=\sum_{t=1}^{M}(Q_{i}c_{j})_{t}.$ (5)
$R_{i}\in\mathbb{R}^{U}$ represents the relevancy score between chunk $i$ and all memory blocks. However, since chunk $i$ should only retrieve memory blocks that came before it temporally, we add a mask to $R_{i}$ and then apply Softmax to obtain the final scores, $\tilde{R}_{i}$, between chunk $i$ and all memory blocks as follows:
$\tilde{R}_{i}$ $=\mathrm{softmax}\left(\frac{1}{\sqrt{d_{\textrm{model}}}}(R_{i}+\mathcal{M}_{i})\right)\in\mathbb{R}^{U}$ (6)
where $\mathcal{M}_{i}$ is a mask of 0 and $\infty$- ∞ constructed to set the scores of future memory blocks (relative to chunk $i$) to $\infty$- ∞. Once all relevancy scores are computed, chunk $i$ simply retrieves the top $k$ memory blocks with the highest $\tilde{R}_{ij}$ scores and concatenates them with the keys and values as in Equations 1 and 2 before computing Attention as in Equation 3.
Compressed Memory Blocks. Next, we discuss how we construct the compressed memory block representations, $c_{j}$. Recently, Landmark Attention [23] considered using “landmark” tokens to obtain compressed representations of memory blocks. We consider using landmark tokens to construct $c_{j}$ in Sec. 8. In SE-Attn, we consider a simpler approach which we found to work well. Namely, for each memory block, $(Q^{\textrm{Mem}}_{j},K^{\textrm{Mem}}_{j},V^{\textrm{Mem}}_{j})$, we perform standard non-causal Attention, $A^{\textrm{Mem}}_{j}=\mathrm{softmax}\left(\frac{Q^{\textrm{Mem}}_{j}(K^{\textrm{Mem}}_{j})^{T}}{\sqrt{d_{\textrm{model}}}}\right)V^{\textrm{Mem}}_{j}\in\mathbb{R}^{S\times d_{\textrm{model}}}$; we consider non-causal Attention since we are interested in a global representation where all tokens within the memory block can attend to each other. Next, we simply compute the mean of these weighted memory tokens as the compressed representation as:
$c_{j}$ $=\frac{1}{S}\sum_{t=1}^{S}(A^{\textrm{Mem}}_{j})_{t}\in\mathbb{R}^{d_{\textrm{model}}}$ (7)
where $(A^{\textrm{Mem}}_{j})_{t}$ denotes the $t$th row of $A^{\textrm{Mem}}_{j}$.
3.4 Training SE-Attn with LoRA
In this paper, we fine-tune models pre-trained with standard Attention; however, we fine-tune them with SE-Attn, which modifies the pre-trained Attention mechanism by introducing a retrieval mechanism. SE-Attn repurposes the model’s Attention parameters $(W_{Q},W_{K},W_{V},W_{o})$ to perform retrieval. In order to efficiently train the model to learn to use SE-Attn, we use a variant of LoRA. Recently, [3] introduced LoRA+, a variant of LoRA [17] designed to fine-tune Transformer models on long contexts. LoRA fine-tunes Attention parameters with low rank adapter matrices. LoRA+ differs from LoRA by also training embedding and normalization layers.
Recently, [10] found that pure SSMs can be fine-tuned by training the SSM projection layers with LoRA. Since we fine-tune on long contexts, we prioritize efficient training, and consequently apply LoRA only to the Attention layers of our hybrid models. However, it is common for SSM layers to include a 1D convolution layer after their initial projection in order to perform sequence mixing [38, 12, 22, 7, 8]. The 1D convolution parameters constitute a very small portion of the model parameters ($\sim$0.7for Mamba-2-Hybrid 2.7B), but as discussed further in Sec. 4.6, we found that training these 1D convolution layers in conjunction with LoRA+ improved our models’ performance on long-context tasks. We refer to this augmented LoRA+ variation as HyLoRA and use it to fine-tune all of our hybrid models.
4 Experiments
4.1 Experimental Setup
Models. We enhance the recall capabilities of pre-trained hybrid SSM models by fine-tuning them with different Attention layers on spans of tokens longer than the ones used for pre-training. We consider Mamba-2-Hybrid 2.7B [^3] We use the Mamba-2-Hybrid weights available on HuggingFace: https://huggingface.co/state-spaces/mamba2attn-2.7b. [34] as our representative SSM hybrid model. We also explore expanding the span of Transformer models, and consider Llama1 7B [31] as our representative model. Both pre-trained model checkpoints were obtained from HuggingFace and were pre-trained with a context size of 2048. Throughout our experiments, “Non-fine-tuned” refers to the pre-trained model with no fine-tuning.
Datasets. We fine-tune models using a common language dataset and provide additional results when fine-tuning on PG-19 [26] and a mixture of language and code dataset in Sec. 7.
Baselines. For Mamba-2-Hybrid, we compare SE-Attn to three Attention variants: standard full Attention (“Full-Attn”) [33], Sliding Window Attention (“SW-Attn”) [1], and Shifted Sparse Attention ($S^{2}$-Attn) [3]. The Full-Attn baseline serves as the paragon, as the other methods aim to approximate it. For Llama1, due to computational constraints, we only compare against SW-Attn and $S^{2}$-Attn. All Attention implementations are based on FlashAttention-2 [5].
Training Procedure. We adopt the training recipe used in [3] to fine-tune our models, with the exception of using HyLoRA for Mamba-2-Hybrid. Additional training details are provided in Sec. 9.
Evaluation Metrics. We assess the performance of our models across a range of common benchmarks. To assess the predictive capabilities of our fine-tuned models on long sequence lengths, we measure perplexity (PPL) on the PG-19 validation set. To assess performance on more real-world language tasks in the short and long-context settings, we evaluate our models on various LM Evaluation Harness tasks [11]. To measure our models’ performance on long-context tasks, we evaluate on the challenging RULER [16] benchmark. We consider eleven RULER tasks, encompassing needle-in-a-haystack, variable-hopping, and aggregation tasks; we aggregate these metrics into five groups as explained in Sec. 11.4.
4.2 Evaluating with Full Attention
Similar to [3], we utilize SE-Attn for efficient fine-tuning and revert to using Full-Attn during evaluation. Due to KV-caching, the complexity of Full-Attn scales linearly during evaluation, so an efficient Attention layer in this setting is not as crucial as in training. Nevertheless, for the sake of completeness, we begin by evaluating our models with the same Attention layer used during fine-tuning (i.e., we fine-tune and deploy our Hybrid model with SE-Attn). In Tab. 1, we experiment with models that use Full-Attn, SW-Attn, $S^{2}$-Attn, and SE-Attn; then, we evaluate perplexity on the PG-19 dataset. All models were fine-tuned with a context size of 8192. We observe that SW-Attn is best at preserving the perplexity on context sizes up to 32×32$\times$32 × larger than the one used for pre-training. SE-Attn is also able to maintain a lower perplexity across longer context sizes, while $S^{2}$-Attn deteriorates much more quickly.
Attention Evaluation Context Size (PPL $\downarrow$)
8192 16384 32768 65536
Non-fine-tuned 14.99 19.35 26.37 34.51
Full-Attn 10.28 10.39 11.14 12.38
SW-Attn 10.33 10.22 10.16 10.16
$S^{2}$-Attn 10.73 10.76 11.85 13.72
SE-Attn 10.47 10.70 10.91 10.96
Table 1: Mamba-2-Hybrid fine-tuned with SE-Attn or SW-Attn preserves perplexity up to 32×32$\times$32 × the pre-training context size. We fine-tune a Mamba-2-Hyrbid model (pre-trained on a context size of 2048) on a context size of 8192 using various Attention variants. Then, we deploy them on longer context sizes using the same Attention variant used during adaptation.
4.3 Mamba-2-Hybrid
All of our Mamba-2-Hybrid models are fine-tuned with a context size of 8192. For SE-Attn, we use a block size of 32, and a top-$k$ of 8. For SW-Attn, we use a window size of 4096. $S^{2}$-Attn uses the parameters from [3]. When fine-tuning Mamba-2-Hybrid using SE-Attn, we found that applying the same chunk sizes at each layer leads to suboptimal downstream performance, we therefore segment each sample into chunks of variable sizes (picked randomly from $\{2048,4096\}$). We found this prevents the model from ‘overfitting’ the given segmentation of the past which is an hyper-parameter of our method fixed a-priori and not dependent on the actual data content. An ablation on SE-Attn with different chunk sizes is provided in Sec. 4.5.
Perplexity. We provide PG-19 PPL results for Mamba-2-Hybrid in Tab. 2. All fine-tuned models improve upon the non-fine-tuned model. SE-Attn yields the closest performance to the paragon model fine-tuned with Full-Attn, outperforming $S^{2}$-Attn and SW-Attn across all context sizes at or above the fine-tuning size.
LM Harness. Next, we evaluate Mamba-2-Hybrid models across short and long-context tasks in the LM Evaluation Harness suite. Our results are provided in Tab. 2, where we observe that all models perform similarly on short-context tasks—including the non-fine-tuned model—suggesting there is no performance regression when fine-tuning on larger contexts. Furthermore, we observe that fine-tuning with SE-Attn gives the closest performance to fine-tuning with Full-Attn.
Attention Eval Context Size (PG-19 PPL $\downarrow$) Short Context Tasks ($\uparrow$) Long Context Tasks ($\uparrow$) 2048 8192 16384 32768 ARC-E ARC-C Hella. LAMB. PIQA WG Avg. SWDE SQA SNQA Avg. Non-fine-tuned 10.72 14.99 19.35 26.37 69.91 37.97 67.62 69.84 76.06 65.04 64.41 85.60 15.18 3.65 34.81 Full-Attn 10.99 10.28 10.39 11.14 69.53 38.48 67.30 68.93 75.08 64.40 63.95 85.24 26.99 19.75 43.99 SW-Attn 10.98 10.80 11.82 13.45 69.82 38.23 67.35 69.18 75.30 63.85 63.95 84.61 24.85 15.41 41.63 $S^{2}$-Attn 10.87 12.89 14.67 16.37 70.12 38.05 67.39 69.84 75.95 64.56 64.32 86.41 17.44 8.53 37.46 SE-Attn 10.99 10.45 11.14 12.64 70.20 38.65 67.15 69.11 75.57 63.93 64.10 85.96 26.70 18.04 43.57
Table 2: Fine-tuning Mamba-2-Hybrid with SE-Attn outperforms fine-tuning with $S^{2}$-Attn and SW-Attn on long-context natural language tasks. We fine-tune Mamba-2-Hybrid with a context size of 8192 using various Attention variants. We evaluate PG-19 validation perplexity (PPL) and observe that fine-tuning with SE-Attn preserves performance at longer contexts better than $S^{2}$-Attn and SW-Attn. On short-context tasks from the LM Harness suite, all models perform similarly. On long context tasks from the LM Harness suite, SE-Attn outperforms $S^{2}$-Attn and SW-Attn. Abbreviations: Hella=HellaSwag, LAMB=LAMBADA, WG=WinoGrande, SQA=ScrollsQAsper, SNQA=ScorllsNarrativeQA.
RULER. Next, we assess the performance of our models on long-context tasks using the RULER benchmark. We plot the average accuracy across eleven RULER tasks in Figure 2(a) (see Sec. 11.4 for details on how we aggregate these metrics). We observe that fine-tuning with SE-Attn yields a performance similar to fine-tuning with Full-Attn, and outperforms $S^{2}$-Attn and SW-Attn. We provide additional RULER metrics in Figure 8, where we see a substantial improvement on the variable tracking (VT) task, which may be attributed to SE-Attn’s retrieval during fine-tuning.
(a) Mamba-2-Hybrid
(b) Llama1
Figure 2: Fine-tuning with SE-Attn outperforms SW-Attn and $S^{2}$-Attn on the RULER benchmark when applied to Mamba-2-Hybrid and Llama1. We fine-tune Mamba-2-Hybrid with a context size of 8192, and Llama1 with 16384 using various Attention variants. We average over eleven RULER tasks, as explained in Sec. 11.4. Fine-tuning with SE-Attn consistently outperforms SW-Attn and $S^{2}$-Attn even when evaluating on context sizes beyond the fine-tuning size.
4.4 Llama1 7B
When fine-tuning Llama 7B models, we found that using a fixed chunk size of $M=4096$ gave better downstream performance. All of our Llama models are fine-tuned with a context size of 16384. For SE-Attn, we use block size of 32, and a top-$k$ of 8. For SW-Attn, we use a window size of 4096, and $S^{2}$-Attn uses the default parameters in [3]. All Attention variants use the same RoPE [29] scaling as in [3].
Perplexity. Fine-tuning Llama with SE-Attn leads to better generalization on context sizes smaller and larger than the one used for fine-tuning. In Tab. 3, we observe that fine-tuning with SE-Attn preserves the performance of the non-fine-tuned model at smaller context sizes, and offers greater generalization to larger context sizes than $S^{2}$-Attn and SW-Attn, as measured on PG-19 validation perplexity.
LM Harness. Fine-tuning Llama with SE-Attn yields a stronger performance on long-context tasks on the LM Evaluation Harness benchmark. As shown in Tab. 3, fine-tuning with SW-Attn, $S^{2}$-Attn, and SE-Attn all improve upon the non-fine-tuned model on shorter context tasks and perform similarly. However, SE-Attn gives a greater performance on long context tasks.
Attention Eval Context Size (PG-19 PPL $\downarrow$) Short Context Tasks ($\uparrow$) Long Context Tasks ($\uparrow$) 2048 8192 16384 32768 ARC-E ARC-C Hella. LAMB. PIQA WG Avg. SWDE SQA SNQA Avg. Non-fine-tuned 8.63 17.62 105.78 244.32 73.19 41.38 66.64 54.38 76.28 62.51 62.40 38.52 20.75 12.36 23.88 SW-Attn 8.80 8.17 8.35 10.32 74.20 43.09 75.00 71.96 77.42 67.48 68.19 82.81 24.41 21.06 42.76 $S^{2}$-Attn 9.32 8.64 8.58 10.42 74.20 42.58 73.84 69.94 77.80 67.72 67.68 80.56 23.36 19.65 41.19 SE-Attn 8.76 8.13 8.01 9.28 75.04 44.37 75.02 71.03 78.02 67.17 68.44 84.79 24.59 21.36 43.58
Table 3: Fine-tuning Llama1 with SE-Attn outperforms fine-tuning with $S^{2}$-Attn and SW-Attn on natural language tasks. We fine-tune Llama1 with a context size of 16384 using various Attention variants. Similar to applying SE-Attn to Mamba-2-Hybrid, here we again observe that fine-tuning Llama with SE-Attn improves upon SW-Attn and $S^{2}$-Attn .
RULER. Fine-tuning Llama with SE-Attn produces a model with stronger performance on RULER tasks than fine-tuning with SW-Attn and $S^{2}$-Attn, as illustrated in Figure 2(b). On average, SW-Attn and $S^{2}$-Attn perform similarly, however, fine-tuning with SE-Attn improves performance by $\sim$5%. Interestingly, despite all models having a similar PPL up to a context size of 16k as shown in Tab. 3, we observe a substantial difference between the RULER performance of models fine-tuned with our SE-Attn and those fine-tuned with SW-Attn and $S^{2}$-Attn . We expand on this observation for Hybrid models in Sec. 6.
4.5 Ablations on Mamba-2-Hybrid
In this section, we ablate over some of the design choices of SE-Attn on Mamba-2-Hybrid. For this analysis, we consider the RULER benchmark, as it is a strong indicator of performance on long-context tasks. As our metric, we use the average of the eleven RULER tasks defined in Sec. 11.4.
Does retrieval during training help? In Sec. 3.2, we introduced SE-Attn-NoMem, a variant of SE-Attn where we do not do any retrieval and process chunks independently. Naturally, we do not expect this to do well due to the lack of shared information across chunks. We confirm this in Figure 3(a), where we observe that SE-Attn-NoMem achieves a much lower performance than SE-Attn with retrieval. Furthermore, we also fine-tune with SE-Attn-Random, a variant of SE-Attn where we retrieve random memory blocks from the past. We observe that this improves upon SE-Attn-NoMem, indicating that retrieving some information from the past improves upon no information. However, SE-Attn-Random does not do as well as SE-Attn, which decides which blocks to retrieve based on relevancy.
Chunk size. For Mamba-2-Hybrid, we found that using a random chunk size during each forward pass improved upon having a fixed chunk size. In Figure 3(b), we compare the performance of SE-Attn, which chooses a random chunk size in $\{2048,4096\}$ during each forward call, to SE-Attn (2048M) and SE-Attn (4096M), which use a fixed chunk size of 2048 and 4096, respectively. We observe that SE-Attn with a fixed chunk size of 4096 improves upon SE-Attn with a fixed chunk size of 2048, but SE-Attn outperforms both. This suggests that training with random chunk sizes may have a regularizing effect, making the model more robust to changes in the input context size.
Figure 3: SE-Attn ablations on Mamba-2-Hybrid. We plot the average of eleven RULER tasks as described in Sec. 11.4. (a): SE-Attn-NoMem processes chunks of tokens without retrieval, while SE-Attn-Random populates its expansion span by retrieving random memory blocks for each chunk. We observe that our Attention-based memory retrieval (SE-Attn) gives the strongest performance. (b): Using SE-Attn with a chunk size chosen randomly from $\{2048,4096\}$ acts as a regularizer and outperforms SE-Attn with fixed chunk sizes of 2048 and 4096. (c): SE-Attn with larger memory blocks (i.e., more tokens per memory block) with a smaller top-$k$ tends to do better than smaller block sizes with a larger top-$k$. (d): An expansion span consisting of 256 total tokens (8 memory blocks with 32 tokens in each) gives the strongest performance. 32S/32k is omitted due to memory constraints.
Block size and top-$k$. For a fixed expansion span size, we might expect that retrieving memory blocks at a finer granularity should perform better than retrieving a smaller number of large blocks; this is because retrieving a larger number of small blocks is as flexible as retrieving a small number of large blocks. However, as illustrated in Figure 3(c), we found that fine-tuning SE-Attn with larger block sizes and a smaller top-k gave the best performance. This is possibly due to the increased complexity of the retrieval task, which gets more difficult as the number of possible blocks to retrieve increases. As we are fine-tuning with LoRA, we hypothesize that the model’s capacity may not be sufficient to learn to retrieve effectively in this setting. As illustrated in Figure 3(d), we found that an expansion span populated with 256 retrieved tokens (32 memory blocks with 8 tokens in each) works best.
4.6 HyLoRA for Hybrid Models
We fine-tune our hybrid models using HyLoRA, which builds upon LoRA+ [3] by also training 1D convolution layers. As illustrated in Figure 4, when evaluated on long-context RULER tasks, fine-tuning SE-Attn models with HyLoRA gives the strongest performance. See Sec. 6 for additional results when fine-tuning with Full-Attn, and the effect of the LoRA rank on downstream performance.
Figure 4: HyLoRA outperforms LoRA and LoRA+ on Hybrid models. We fine-tune Mamba-2-Hybrid with SE-Attn using LoRA, LoRA+, and HyLoRA. We find that LoRA and LoRA+ perform sub-optimally. HyLoRA augments LoRA+ by additionally training the 1D convolution layers and yields strong performance.
4.7 Empirical Runtime Analysis
SE-Attn is much faster than Full-Attn and $S^{2}$-Attn on large contexts. In Figure 5, we profile different Attention layers on various context sizes and see that SE-Attn is much faster than $S^{2}$-Attn and Full-Attn. Moreover, we see that the runtime of SE-Attn is similar to that of SW-Attn, indicating that our retrieval overhead is minimal despite the much longer Attention span. See Sec. 10 for a more detailed runtime analysis of SE-Attn.
Figure 5: SE-Attn is faster than $S^{2}$-Attn and Full-Attn, especially on long contexts. The retrieval overhead of SE-Attn is minimal, with runtime similar to SW-Attn (which has a limited pre-determined Attention span).
5 Conclusion
We close with the limitations of our method. Although we have conducted experiments at a relatively large model scale (2.7B for Hybrid models and 7B for Transformers) and long contexts (up to 32k), testing our method for efficiently adapting larger models on longer contexts is paramount. Furthermore, while we utilize datasets that require modeling of long-range dependencies, we found that perplexity-based tasks do not faithfully measure models’ capabilities to handle long contexts, and instead tasks like RULER provide better signals of long-context capabilities. However, RULER is mostly a synthetic dataset and does not cover more nuanced tasks that require reasoning over long documents. Validating our method on more complex long-range benchmarks is a promising area for future work.
Supplementary Material
6 LoRA for Hybrid Models and the Pitfalls of Perplexity
In this section, we conduct a more thorough ablation study on the effect of training Mamba-2-Hybrid with different variants of LoRA.
Fine-tuning with Full-Attn. We first expand upon Figure 4, where we showed that fine-tuning Mamba-2-Hybrid with SE-Attn using HyLoRA produced a stronger model than training with LoRA and LoRA+. In Figure 6, we provide a similar plot, but using Full-Attn during fine-tuning rather than SE-Attn. We observe a similar trend, where HyLoRA produces a stronger model than LoRA and LoRA+. In the remaining analyses in this section, we provide results for fine-tuning with SE-Attn, as we observed similar trends when fine-tuning with Full-Attn.
Figure 6: HyLoRA outperforms LoRA and LoRA+ on Hybrid models. We fine-tune Mamba-2-Hybrid with Full-Attn using LoRA, LoRA+, and HyLoRA. We find that LoRA and LoRA+ perform sub-optimally compared to HyLoRA which also trains 1D convolution layers.
Effect of LoRA rank. We next explore how the rank used for LoRA fine-tuning affects downstream performance on the RULER task. Given its stronger performance, we focus on our HyLoRA. In our previous Mamba-2-Hybrid experiments, we used a LoRA rank $(r)$ of 32 and an $\alpha$ of 64 (for our Transformer experiments, we used $r=8$ and $\alpha=16$ as in [3]). In Figure 7, we plot the average RULER results for Mamba-2-Hybrid models fine-tuned with HyLoRA using different ranks (we scale $\alpha$ to maintain the ratio $\alpha/r=2$). Here, we observe that training with a larger rank improves downstream performance, and we start to see some saturation around $r=64$.
In Tab. 4, we provide PG-19 perplexity results for Mamba-2-Hybrid trained with different LoRA variants. All models are fine-tuned with a context size of 8192, and evaluated with multiple context sizes. For a given context size, we do not see a substantial difference in the perplexity. This is similar to the observation in [3]. However, as discussed above, the rank can have a significant effect on downstream long-context tasks that require strong recall capabilities, as in RULER. Hence, while perplexity results may be promising, they are not necessarily indicative of performance on more complex long-context tasks.
Based on the analyses in this section, we conclude the following:
  • Fine-tuning the 1D convolution layers, as we do in our HyLoRA, significantly improves performance on downstream long-context tasks that require retrieval, such as RULER.
  • Fine-tuning with larger LoRA ranks improves performance up to a certain point—64 for our experiments.
  • Perplexity may not be the most faithful metric for assessing performance on long-context downstream tasks. Instead, researchers should consider evaluating on more complex tasks, such as those in the RULER benchmark.
Figure 7: Fine-tuning with a larger LoRA rank using HyLoRA improves performance on the RULER benchmark. We fine-tune Mamba-2-Hybrid using HyLoRA with different LoRA ranks (we maintain a LoRA rank to alpha ratio of 2). We observe that fine-tuning with a larger rank produces stronger downstream results on RULER, with some saturation with a rank of 64.
Sequence Length LoRA Method (Rank=32) LoRA Rank (Method=HyLoRA) LoRA LoRA+ HyLoRA 8 16 32 64 2048 10.77 10.98 10.99 11.00 11.00 10.99 10.99 8192 10.29 10.49 10.45 10.45 10.46 10.45 10.44 16384 10.91 11.18 11.14 11.03 11.10 11.14 11.14 32768 12.34 12.66 12.64 12.37 12.53 12.64 12.64 65536 13.99 14.36 14.26 13.84 14.08 14.26 14.28
Table 4: Mamba-2-Hybrid LoRA Ablations. We fine-tune Mamba-2-Hybrid with a context size of 8192 with SE-Attn using different LoRA variants. We consider LoRA, LoRA+, and HyLoRA (ours) and evaluate perplexity on the PG-19 dataset. We observe that all LoRA variants yield similar perplexity results. However, as depicted in Figure 4 and Figure 7, different LoRA variants yield substantially different performances on more complex tasks.
7 Additional Fine-Tuning Results
In this section, we provide additional benchmarks on the Mamba-2-Hybrid model. We begin by expanding upon Figure 2(a) with additional RULER metrics in Figure 8. Here, we observe that Mamba-2-Hybrid fine-tuned with SE-Attn consistently outperforms SW-Attn and $S^{2}$-Attn on all RULER tasks across a broad range of context sizes—even beyond the 8192 fine-tuning context size.
Figure 8: Mamba-2-Hybrid RULER benchmark fine-tuned on natural language. We fine-tune Mamba-2-Hybrid with a context size of 8192 using multiple Attention variants and evaluate on the RULER benchmark (please see Sec. 11.4 for definitions of each metric). On average, fine-tuning with SE-Attn yields a performance close to the paragon model fine-tuned with Full-Attn . We observe the biggest gain in needle-in-a-haystack (NIAH) tasks, as well as variable tracking (VT), all of which require strong recall capabilities enabled by our SE-Attn.
7.1 Fine-Tuning on Natural Language + Code
Next, we consider fine-tuning on a dataset that augments natural language with a code dataset. In particular, we construct a dataset consisting of 70% natural language, and 30% C++ code extracted from the Lots of Code [39] dataset. We provide PG-19 validation PPL results, as well as results on tasks from the LM Evaluation Harness suite in Tab. 5. Similar to fine-tuning on natural language only (as in Tab. 2), we observe that fine-tuning with SE-Attn yields the strongest performance on downstream tasks. We provide performance on RULER in Figure 9. We again observe that fine-tuning with SE-Attn yields the strongest perfromance compared to other efficient Attention layers. Moreover, compared to fine-tuning only on natural language, here we observe a greater improvement on the Variable Tracking task. As explained in Sec. 11.4, this task requires the model to keep track of the values of variables that are defined and overridden throughout the input context, and must then return the variables equal to some value. This requires strong recall capabilities, which fine-tuning with SE-Attn enables, and is amplified by the use of code data in the fine-tuning data mix.
Attention Eval Context Size (PG-19 PPL $\downarrow$) Short Context Tasks ($\uparrow$) Long Context Tasks ($\uparrow$) 2048 8192 16384 32768 ARC-E ARC-C Hella. LAMB. PIQA WG Avg. SWDE SQA SNQA Avg. Non-fine-tuned 10.72 14.99 19.35 26.37 69.91 37.97 67.62 69.84 76.06 65.04 64.41 85.60 15.18 3.65 34.81 Full-Attn 10.94 10.23 10.36 11.09 70.12 38.14 67.40 69.34 75.08 64.56 64.11 84.88 25.99 19.61 43.49 SW-Attn 10.94 10.76 11.81 13.43 69.70 38.82 67.51 69.38 75.41 64.88 64.28 84.52 24.44 15.30 41.42 $S^{2}$-Attn 10.86 12.89 14.67 16.36 69.95 37.88 67.45 69.94 76.01 64.72 64.32 86.50 16.65 8.61 37.25 SE-Attn 10.95 10.41 11.07 12.50 70.45 38.91 67.39 69.07 75.08 64.96 64.31 85.24 26.14 18.08 43.15
Table 5: Fine-tuning Mamba-2-Hybrid with SE-Attn on a natural language + code dataset outperforms fine-tuning with $S^{2}$-Attn and SW-Attn on natural language tasks. We fine-tune Mamba-2-Hybrid with a context size of 8192 using various Attention variants on a dataset that consists of 70% natural language and 30% code. We evaluate PG-19 validation perplexity (PPL) and observe that fine-tuning with SE-Attn yields better perplexity scores than $S^{2}$-Attn and SW-Attn. On short-context tasks from the LM Harness suite, all models perform similarly. On long context tasks from the LM Harness suite, SE-Attn outperforms $S^{2}$-Attn and SW-Attn.
Figure 9: Mamba-2-Hybrid RULER benchmark fine-tuned on natural language + code. We fine-tune Mamba-2-Hybrid with different Attention layers on a dataset that consists of 70% natural language and 30% code and evaluate on RULER. Compared to fine-tuning on only natural language (as in Figure 8), we see a substantial improvement on tasks like variable tracking (VT), and needle-in-a-haystack tasks (NIAH), both of which require strong recall capabilities enabled by fine-tuning with SE-Attn.
7.2 Fine-Tuning on PG-19
Next, we consider fine-tuning on the PG-19 [27] dataset. We provide perplexity results on the validation set, along with LM Harness task results in Tab. 6. Compared to fine-tuning on a different natural language dataset, and a natural language + code as in Tab. 2 and Tab. 5, here we obtain lower perplexity scores due to the lack of distribution shift. Interestingly, compared to fine-tuning on the previously mentioned datasets, we observe a slight degradation on LM Harness tasks. Moreover, in Figure 10 we plot RULER performance when fine-tuning on PG-19, and here we also see a decay in performance compared to fine-tuning on the previous datasets. This suggests that the PG-19 dataset may be too far out of distribution for these long-context retrieval task. Nevertheless, we see that fine-tuning with SE-Attn yields the best performance across all of these tasks.
Attention Eval Context Size (PG-19 PPL $\downarrow$) Short Context Tasks ($\uparrow$) Long Context Tasks ($\uparrow$) 2048 8192 16384 32768 ARC-E ARC-C Hella. LAMB. PIQA WG Avg. SWDE SQA SNQA Avg. Non-fine-tuned 10.72 14.99 19.35 26.37 69.91 37.97 67.62 69.84 76.06 65.04 64.41 85.60 15.18 3.65 34.81 Full-Attn 10.73 10.04 10.14 10.91 67.34 37.12 66.80 68.62 74.32 62.67 62.81 84.88 25.35 18.46 42.90 SW-Attn 10.72 10.59 11.64 13.25 66.96 37.54 66.83 68.79 74.54 63.30 62.99 84.79 22.54 14.09 40.48 $S^{2}$-Attn 10.78 12.74 14.42 16.11 69.36 38.14 67.45 69.90 76.12 64.72 64.28 86.50 17.00 7.74 37.08 SE-Attn 10.73 10.22 10.84 12.28 67.42 37.88 66.48 69.22 73.88 61.88 62.80 85.15 23.06 16.46 41.55
Table 6: Fine-tuning Mamba-2-Hybrid with SE-Attn on PG-19 outperforms fine-tuning with $S^{2}$-Attn and SW-Attn on long-context natural language tasks. We fine-tune Mamba-2-Hybrid with a context size of 8192 using various Attention variants on PG-19. We evaluate PG-19 validation perplexity (PPL) and observe that fine-tuning with SE-Attn yields better perplexity scores than $S^{2}$-Attn and SW-Attn. On short-context tasks from the LM Harness suite, $S^{2}$-Attn has the strongest performance. On long context tasks from the LM Harness suite, SE-Attn outperforms $S^{2}$-Attn and SW-Attn.
Figure 10: Mamba-2-Hybrid RULER benchmark fine-tuned on PG-19. We fine-tune Mamba-2-Hybrid with different Attention layers on the PG-19 [27] dataset and then evaluate on RULER. Compared to fine-tuning on other natural language datasets with a greater variety of text as in Figure 8, here we observe a degradation in performance across all models, likely due to a distribution shift in the PG-19 data and the RULER tasks.
7.3 Evaluating with Efficient Attention
In Sec. 4.2, we explained how we fine-tune our models using an efficient Attention mechanism, and then evaluate with full attention. In Tab. 1, we showed that evaluating our models with efficient Attention mechanisms produced strong results on perplexity benchmarks. However, we did not observe similarly positive results on more complex tasks, such as those in RULER. As illustrated in Figure 11, using the same efficient Attention layer used during fine-tuning does not perform as well as evaluating with Full-Attn (we omit results for models fine-tuned with $S^{2}$-Attn and evaluated with $S^{2}$-Attn as the implementation of $S^{2}$-Attn does not support evaluating with it at arbitrary sequence lengths). Hence, our training pipeline involves training a model with an efficient Attention layer, and reverting to Full-Attn during inference. Moreover, we note that perplexity is again a misleading metric in regard to measuring a model’s performance on long context tasks. Though we observed strong performance on perplexity when evaluating with these efficient Attention layers, this did not translate to stronger performance on tasks that require a strong recall capability, such as those in RULER.
Figure 11: Evaluating with efficient Attention mechanisms on RULER does not do as well as evaluating with standard full Attention. We fine-tune Mamba-2-Hybrid with SW-Attn (a) and SE-Attn (b) and then evaluate on the NIAH-Single-1 RULER task using either the same Attention layer used for fine-tuning, or Full-Attn. We observe that fine-tuning with efficient Attention layers (SE-Attn and SW-Attn) and then using Full-Attn during evaluation yields better results. Lines with a circle marker denote models fine-tuned with an efficient Attention mechanism, and then evaluated with the same Attention mechanism; lines with a diamond marker were fine-tuned with an efficient Attention mechanism, but evaluated with Full-Attn. We omit results for evaluating models fine-tuned with $S^{2}$-Attn as this Attention mechanism does not support arbitrary sequence lengths during inference.
8 Retrieval with Landmark Tokens
Landmark Attention [23] was recently introduced as way for Transformer models to process long sequences by inserting ‘landmark’ tokens into the sequence whose representations would then be used as summaries of the blocks of tokens that came before them. At a high level, our approach in SE-Attn is similar. However, we simplify the process of compressing blocks of tokens by forgoing the use of landmark tokens and instead using Attention to summarize them, as described in Sec. 3.3. Moreover, to learn which blocks to retrieve, we do not rely on a complex Grouped Softmax function, and instead use a simple Cross-Attention score to ascertain relevance. In this way, we implement retrieval natively into the model’s architecture.
We consider a variant of SE-Attn, which we refer to as SE-Attn-LM, that is inspired by Landmark Attention. Namely, instead of using our Attention-based compression to construct summaries of memory blocks, we insert a non-learnable ‘landmark’ token into each memory block, and use the cross-attention between this token and the memory tokens as the summary of the memory block. We compare the performance of this variant to our summaries computing using average pooling of Attention scores (see Equation 7) in Figure 12. Here, we see that SE-Attn yields better performance. We suspect this is because using a non-learnable landmark token to summarize memory blocks is too challenging of a task to accomplish using standard LoRA. While full fine-tuning (without LoRA) may improve the performance of SE-Attn-LM, this is beyond the scope of our work as we prioritize efficiency.
Figure 12: Summarizing memory blocks via average pooling of attention yields stronger performance than summarizing them using ‘landmark’ tokens. We fine-tune Mamba-2-Hybrid using our SE-Attn, and SE-Attn-LM, which summarizes memory blocks with a landmark token (similar to [23]) instead of using the average of the Self Attention output of memory blocks, as in SE-Attn. We find that our simpler SE-Attn produces a stronger model, likely due to the easier training task, which does not require adapting the model to leverage landmark tokens for compression.
9 Training Details
Our fine-tuning recipe largely follows [3], with the exception of using a larger learning rate for SE-Attn, SW-Attn, and Full-Attn models ($2\times 10^{-4}$ vs. the default $2\times 10^{-5}$ for $S^{2}$). We found that using a larger learning rate for $S^{2}$ did not improve its performance, as shown in Figure 13, so we used the default $2\times 10^{-5}$ learning rate for all $S^{2}$ fine-tuning experiments.
All of our experiments are conducted on a single 8xA100 node. We fine-tune for 1000 steps with a total of 0.5B tokens. We fine-tune Mamba-2-H with 8192 tokens with 8 accumulation steps and a per-device batch size of 1. We fine-tune Llama1 with 16384 tokens with 4 accumulation steps and a per-device batch size of 1. We use FlashAttention-2 [5] and DeepSpeed Stage 2 [28].
Figure 13: A larger learning rate does not improve the performance of $S^{2}$. Here we fine-tune a Mamba-2-Hybrid model using $S^{2}$-Attn with two different learning rates: $2\times 10^{-4}$ and $2\times 10^{-5}$. The learning rate used in [3] is $2\times 10^{-5}$, which we found to work well. For all other Attention layers, we found $2\times 10^{-4}$ to offer a slight improvement over $2\times 10^{-5}$.
10 Runtime Analysis
The runtime of SE-Attn is lower than Full-Attn and $S^{2}$-Attn on large contexts. For an input with sequence length $L$, SE-Attn first constructs $U=\frac{L}{S}$ memory blocks of size $S$. Attention is applied on each. Thus, the complexity of this operation is $O(d_{\textrm{model}}US^{2})=O(d_{\textrm{model}}LS)$. Next, the input sequence is split into $T=\frac{L}{M}$ chunks of size $M$. Cross-Attention is then applied between each chunk’s query and each memory block’s compressed representation; the cost of this is $O(d_{\textrm{model}}MTU)=O(d_{\textrm{model}}\frac{L^{2}}{S})$. Cross-Attention is then applied between each chunk’s query tokens and the concatenations of the chunk’s key tokens and memory tokens. Since SE-Attn retrieves $K$ blocks with $S$ tokens in each, the cost of this Cross-Attention for each chunk is $O(d_{\textrm{model}}M(SK+M))=O(d_{\textrm{model}}M^{2})$ since $SK<M$. The cost for all chunks is therefore $O(d_{\textrm{model}}TM^{2})=O(d_{\textrm{model}}LM)$. The total cost of SE-Attn is therefore $O(d_{\textrm{model}}LS+d_{\textrm{model}}LM+d_{\textrm{model}}\frac{L^{2}}{S})$. For sufficiently large $S$, the runtime of SE-Attn is faster than Full-Attn and $S^{2}$-Attn, especially on large contexts, as shown empirically in Sec. 4.7.
11 RULER Task Definitions
The RULER benchmark [16] consists of four different task categories: retrieval, multi-hop tracing, aggregation, and question answering. In this paper, we focus only on the retrieval, multi-hop tracing, and aggregation tasks (we provide question answering results on the LM Evaluation Harness benchmark). These three categories span eleven different tasks as explained below.
11.1 Needle-in-a-Haystack (NIAH) Tasks
RULER consists of 8 different NIAH tasks. These tasks embed “needles” in a string of noise. These needles are typically key-value pairs, and the goal is to return the value of a key. These tasks are characterized by six parameters:
  • type_haystack (TH): This specifies the type of noise to embed the key in. The choices are “repeat” which constructs noise as in [23], “essay” which will use sentences from the Paul Graham essays [20], or “needle” in which case each sentence will define a new key-value pair.
  • type_needle_k (TK): This specifies the type of the needle’s key. The options are “words”, in which case the key is a word (in the form of adjective-noun, e.g., spiritual-oven), or “uuids” in which case the key is a UUID.
  • type_needle_v (TV): This specifies the type of the needle’s value. It can either be “numbers” in which case the value is a 7-digit number, or it can be “uuids” in which case the value is a UUID.
  • num_needle_k (NK): This specifies the number of key-value pairs to embed in the haystack.
  • num_needle_v (NV): This specifies how many different values a key is assigned. If greater than 1, the goal is output all the values of they key.
  • num_needle_q (NQ): This specifies the number of different keys the model must return the value for.
NIAH Task TH TK TV NK NV NQ
Single 1 repeat words numbers 1 1 1
Single 2 essay words numbers 1 1 1
Single 3 essay words uuids 1 1 1
Multikey 1 essay words numbers 4 1 1
Multikey 2 needle words numbers 1 1 1
Multikey 3 needle uuids uuids 1 1 1
Multivalue essay words numbers 1 4 1
Multiquery essay words numbers 1 1 4
Table 7: RULER NIAH definitions. The ‘Needle-in-a-Haystack’ (NIAH) tasks in the RULER benchmarks are defined by 6 parameters which modulate the difficulty of the tasks. We consider 8 different NIAH tasks as defined above (these are the default NIAH tasks in the RULER library).
11.2 Multi-hop Tracing Tasks
RULER considers a “variable tracking” task that is a form of coreference resolution. In this task, a sequence of variables are defined throughout noisy text as in [23]. New variables are defined as previous ones, and a final value is assigned to a particular variable. The goal is to be able to trace back which variables have also been assigned the final value, i.e., determine which variables refer to the final value. We use the default num_chains=1 and num_hops=4 parameters
11.3 Aggregation Tasks
RULER considers two aggregation tasks, common words extraction (CWE), and frequent words extraction (FWE). In CWE, the context consists of list of words, and the goal is to return the most common words. We use the default parameters freq_cw=30, freq_ucw=3, and num_cw=10. In FWE, the context consists of random word strings, and the goal is to return the ones that appear the most frequently. We use the default alpha=2 parameter for this.
11.4 Aggregating RULER Tasks
We aggregate the eleven RULER tasks above into six groups:
  1. 1.
    NIAH-S: NIAH Single 1, NIAH Single 2, NIAH Single 3.
  2. 2.
    NIAH-M: NIAH Multikey 1, NIAH Multikey 2, NIAH Multikey 3.
  3. 3.
    NIAH-M-QV: NIAH Multivalue, NIAH Multiquery.
  4. 4.
    VT: Variable Tracking.
  5. 5.
    CF-WE: Common Words Extraction (CWE) and Frequent Words Extraction (FWE).
  6. 6.
    Average: The average of all eleven tasks above.